Please show and tell me how to convert the fraction 341/120 into a usable fraction such as 2 3/8. If you get 2.84167 as the answer, what is that in fraction form such as inches?
Answer by Mathtut(3670)

you divide 342 by 120...it divides in evenly, 2 times and you have 102 left over as a remainder so you write this in the form of 2 102/120. Now you reduce the fractional part as low as it can reduce 102/120=51/60=17/20
:
so we have a final mixed fraction of 2 17/20
:
when you have an answer of 2.84167 and wish to convert to a mixed fraction. first you take the fractional part and if you want to be precise you do not round off. Write it with all the numbers that are to the right of the decimal point. So in this case it would be 84167/100000. You then reduce this as far as possible. If you want just be close and approximate,
:
then you would round to say the nearest 100th place.
:
84/100= 42/50= 21/25.....so that would be 2 21/25
:
hope that helps
: